OpenStudy (anonymous):

Convolution theorem laplace { integral from 0 to t (e^(-tau)cos(tau) d(tau)}

OpenStudy (anonymous):

f(tau) and g(t-tau) how should I manipulate e^(-tau) so that I could extract the function of g?

OpenStudy (anonymous):

still doesn't sound right f(tau) and g(t-tau) f*g= integral 0-->t f(tau)g(t-tau)dtau if f(tau)=cos(tau) and g(t-tau)=e^(-tau) how should i manipulate the function in order to get the funtion of g

OpenStudy (accessdenied):

My first thought was something like this: \( e^{- \tau} = e^{t - \tau - t} = e^{-t} e^{t - \tau} \) Adding and subtracting t.

OpenStudy (accessdenied):

You get an extra e^(-t), which can be pulled outside of the integral because it does not depend on tau. From that though, I think there is a theorem to work with the e^(-t) multiplied to it.

OpenStudy (anonymous):

for this problem, am suppose to transform it back to f*g and take the laplace and get F(s)G(s) though

OpenStudy (anonymous):

and according to the solution manual 1/s Laplace[e^(-t) cost} is the next step am not sure where they pull out the 1/s

OpenStudy (accessdenied):

Strange. I was looking at a table and found a specific identity that gives exactly that. But it did not explain how it was derived in terms of convolutions. Unless it is just calling one function 1? \( \mathcal{L} \left[ f(v) \ dv \right) = \dfrac{F(s)}{s} \) The way I was going about it does not seem to lead there.

OpenStudy (accessdenied):

\( \displaystyle \mathcal{L} \left[ \int_{0}^{t} f(v) \ dv \right] = \dfrac{F(s)}{s} \) Fixed that...

OpenStudy (anonymous):

I suppose...that's the only way I could explain the 1/s

OpenStudy (anonymous):

what exactly is the convolution? lol

OpenStudy (accessdenied):

So, the g(t - tau) = 1, and f(tau) = e^(-tau) cos (tau). I've honestly never seen that one before. o.O

OpenStudy (anonymous):

hmmm, it still hard to picture where that 1/s came from if g(t - tau) = 1, and f(tau) = e^(-tau) cos (tau was the case. I'll just flow with it!

OpenStudy (anonymous):

wait... Laplace{1} =1/s... ahh

OpenStudy (anonymous):

thanks again!

OpenStudy (accessdenied):

Yup, glad to help! I learned something here too.. lol I cannot seem to get the same answer with my original route, either. So I guess that was it.